OpenStudy (anonymous):

y=x2-x-6 y=2x-2 solve. Find coordinates

OpenStudy (shinalcantara):

\[y = x^2 - x - 6\] \[y = 2x - 2\] is this the given?

OpenStudy (anonymous):

yes but these are the options {(2,2)}{(-1,1), (2,4)){(0,1), (1,2)}{(1,3), (6,13)}{(4,6), (-1,-4)}{(5,-1), (1,-5)}

OpenStudy (shinalcantara):

\[y=x^2-x-6\] \[y=2x-2\] Equate the two equations \[x^2 - x - 6 = 2x - 2\] \[x^2 - x -2x - 6 + 2 = 0\] \[x^2 - 3x - 4=0\] \[(x-4)(x+1)=0\] \[x=4; x=-1\] --------------- If x= 4 y = 2x - 2 y = 2(4) - 2 y = 8 - 2 y = 6 (4,6) -------------- If x=-1 y = 2x - 2 y = 2(-1) - 2 y = -2 - 2 y = -4 (-1,-2)